Solve the system by addition.
–3x + 5y = 8
x + 2y = 1

Answer by checkley71(8403)

-3X+5Y=8
X+2Y=1 MULTIPLYING THE SECOND EQUATION BY 3 WE GET
3X+6Y=3NOW WE ADD EQUATION ONE TO THIS EQUATION
-3X+5Y=8
--------
0+11Y=11 OR Y=1 SUBSTIRUTING 1 FOR Y WE NOW GET -3X+5*1=8 OR -3X=3 OR X=-1
PROOF -1=2*1=1 OR -1+2=1 OR 1=1